# VCDiag: Classifying Erroneous Waveforms for Failure Triage Acceleration

### v2

@@ I Introduction
Design functional verification (DFV) is critical for ensuring chips meet specifications, avoiding costly errors, and protecting companies’ reputations. However, traditional methods struggle with modern design complexities. Verification now consumes more engineering resources than design itself, reflected by a rising verification-to-design engineer ratio [1]. Furthermore, verification engineers spend 47% of their time debugging, a 6% increase over four years [2], contributing to 75% of projects missing deadlines and only 14% avoiding a respin due to functional failures [1]. As design complexity grows, innovative debugging tools and methodologies are urgently needed.”
@@ II Previous Work
BugMD [8] detects post-silicon CPU bugs in FabScalar [14] by aligning architectural states with a golden reference and extracting mismatch features for bug signature generation. Using synthetic gate-level bug injection and a random forest classifier, it achieves 70% top-1 and 90% top-3 accuracy across twelve CPU regions. Since post-silicon timing data and gate-level bug injection are unavailable pre-synthesis, we target a similar problem at the RTL level. We adopted FabScalar’s RTL design, applied comparable metrics, and developed a targeted RTL bug generation tool. Our approach matches or exceeds BugMD’s reported results on FabScalar and consistently performs well across a broader set of modern, open-source designs, demonstrating its scalability and applicability.
@@ III Problem Formulation
A Verilog/SystemVerilog module defines the behavior of a hardware block through internal logic and signal interactions. Modules communicate via well-defined interfaces and can be composed hierarchically to build complex digital systems. During simulation, the complete design, referred to as the Design-Under-Test (DUT), is exercised by testbenches that apply inputs and check outputs. When a mismatch is detected, the testbench flags the failing cycle and logs diagnostic details. DFV engineers then isolate the failing test case, review the logs, and inspect relevant waveforms against the design specification to identify anomalies and locate the root cause.
For our analysis, the simulation halts upon detecting the first error, and the resulting VCD waveform captures DUT behavior at every simulation tick from initialization to failure. Each waveform is labeled with the module responsible for the failure and serves as a behavioral signature for training ML models as shown in Figure 1. A separate model is trained and tested for each design to learn its unique behavior. We formulate this as a multivariate time-series classification (MTSC) problem, where a waveform is represented as $X\in\mathbb{R}^{T\times d}$, with $T$ time steps and $d$ signal measurements per step. The objective is to learn a model $f:\mathbb{R}^{T\times d}\to\{1,2,\dots,M\}$ that maps a failing waveform to the corresponding module label $Y$.
@@ IV Proposed Framework
Figure 2: VCDiag framework
This section describes our methodology for extracting structured features from raw VCDs, collecting labeled waveform data from regression runs, and generating synthetic bugs to support model training. We then present the ML models evaluated and explain the rationale behind their use for waveform classification.
@@ IV-B Parallel simulation, extraction, and processing
    Statistical compression: We compress each signal into one CSV row using statistical metrics like mean, standard deviation, and quantiles. This shrinks data size but increases features. The SummaryTransformer [16] function generates one CSV per bug scenario, with rows as waveforms and $d\times n$ feature columns, where $n$ is the statistical metric count.
@@ IV-C LLM-based synthetic bug generator
A major challenge in evaluating our framework is the scarcity of publicly available RTL-level bug datasets. To address this, inspired by recent advances in Large Language Model (LLM) in source code mutation [17, 18], we developed an LLM-driven bug generation tool for Verilog/SystemVerilog.
@@ V Experimental Setup
Simulator Synopsys VCS© Cadence Xcelium©
@@ VI-B Resource utilization results
Although all experiments were capped at 2000 ticks, the results in Figure 8 suggest that comparable performance can be achieved with as few as 200 ticks, indicating a substantial opportunity for improving efficiency.
@@ VI-C Adaptation
Integrating VCDiag with OpenTitan RomCtrl initially took about three days, as we navigated testbench behavior and post-simulation data handling. After initial integration, extending to additional OpenTitan designs required only hours due to shared UVM infrastructure. For FabScalar, the setup required two days to adapt the framework to its distinct flow. By open-sourcing VCDiag, we aim to foster broader adoption and evolution of our framework in both research and industry.
@@ VI-E Future Work
VCDiag represents a significant step toward developing a production-ready, ML-based failure triage system, opening several promising avenues for future enhancements. First, leveraging standardized logs with support from LLMs and automated log templates will greatly improve log clustering, thereby enhancing coarse-grain debugging by efficiently grouping related failures. Additionally, extending VCDiag to support scenarios with multiple concurrent faulty modules through an iterative ’onion-peeling’ methodology will significantly broaden its practical applicability. This iterative refinement, however, would require developing strategies to accommodate intermediate design modifications seamlessly. Finally, focusing on improving the efficiency of regression reseeds, either through optimized simulation strategies or intelligent prioritization of test scenarios, will substantially accelerate the generation of comprehensive bug datasets. Addressing these aspects will collectively elevate VCDiag towards a robust, industry-ready ML-powered debugging framework.

### v3

@@ Abstract
publicationid: pubid: 979-8-3315-3762-3/25/\$31.00 © 2025 IEEE
Figure 1: VCDiag framework
@@ I Introduction
Design functional verification (DFV) is critical for ensuring chips meet specifications, avoiding costly errors, and protecting companies’ reputations. However, traditional methods struggle with modern design complexities. Verification now consumes more engineering resources than design itself, reflected by a rising verification-to-design engineer ratio [1]. Nearly 75% of projects miss deadlines and only 14% avoid a respin due to functional failures [1]. Furthermore, verification engineers spend 47% of their time debugging [1]. As design complexity grows, innovative debugging tools and methodologies are urgently needed to drive efficiency improvements in verification.
@@ II Previous Work
BugMD [7] detects post-silicon CPU bugs in FabScalar [13] by aligning architectural states with a golden reference and extracting mismatch features for bug signature generation. Using synthetic gate-level bug injection and a random forest classifier, it achieves 70% top-1 and 90% top-3 accuracy across twelve CPU regions. Since post-silicon timing data and gate-level bug injection are unavailable pre-synthesis, we target a similar problem at the RTL level. We adopted FabScalar’s RTL design, applied comparable metrics, and developed a targeted RTL bug generation tool. Our approach exceeds BugMD’s reported results on FabScalar and consistently performs well across a broader set of modern, open-source designs, demonstrating its scalability and applicability.
@@ III Problem Formulation
A Verilog/SystemVerilog module defines the behavior of a hardware block through internal logic and signal interactions. Modules communicate via well-defined interfaces and can be composed hierarchically to build complex digital systems. During simulation, the complete design, referred to as the Design-Under-Test (DUT), is exercised by testbenches that apply inputs and check outputs against a golden reference. When a mismatch is detected, the testbench flags the failing cycle and logs diagnostic details. DFV engineers then isolate the failing test case, review the logs, and inspect relevant waveforms against the design specification to identify anomalies and locate the root cause.
For our analysis, the simulation halts upon detecting the first error, and the resulting VCD (waveform) captures DUT behavior at every simulation tick from initialization to failure. Each VCD is labeled with the module responsible for the failure and serves as a behavioral signature for training ML models as shown in Figure 2. A separate model is trained and tested for each design to learn its unique behavior. We formulate this as a multivariate time-series classification (MTSC) problem, where a waveform VCD is represented as $X\in\mathbb{R}^{T\times d}$, with $T$ time steps and $d$ signal measurements per step. The objective is to learn a model $f:\mathbb{R}^{T\times d}\to\{1,2,\dots,M\}$ that maps a failing waveform to the corresponding module label $Y$.
@@ IV Proposed Framework
This section describes the main components of VCDiag in practical execution order, as shown in Figure 1, and concludes with a discussion of the evaluated machine learning models and the rationale for their selection in failing waveform classification.
@@ IV-B Parallel simulation, extraction, and processing
    Statistical compression: We compress each signal into one CSV row using statistical metrics like mean, standard deviation, and quantiles. This shrinks the data size but increases the features. The SummaryTransformer [15] function generates one CSV per bug scenario, with rows as waveforms and $d\times n$ feature columns, where $n$ is the statistical metric count.
@@ IV-C LLM-based synthetic bug generator
A major challenge in evaluating our framework is the scarcity of publicly available RTL-level bug datasets. To address this, inspired by recent advances in Large Language Model (LLM) in source code mutation [16, 17], we developed an LLM-driven bug generation tool for Verilog/SystemVerilog [18].
@@ V Experimental Setup
Simulator Synopsys VCS [21] Cadence Xcelium [22]
@@ VI-B Resource utilization results
Despite the 2000-tick cap, Figure 8 shows comparable accuracy at 200 ticks, a major efficiency gain opportunity. This is especially evident in FabScalar, where the marginal difference is likely due to many failures being timeouts, allowing the model to identify the faulty module with a small number of ticks.
@@ VI-C Adaptation
Initial OpenTitan RomCtrl integration took three days due to testbench and post-simulation handling, but extending to additional OpenTitan designs required only hours due to shared UVM infrastructure. For FabScalar, the setup required two days to adapt the framework to its distinct flow. By open-sourcing VCDiag, we aim to foster broader adoption and evolution of our framework in both research and industry.
@@ VI-E Future Work
VCDiag represents a significant step toward developing a production-ready, ML-based failure triage system, opening several promising avenues for future enhancements. First, leveraging standardized logs with support from LLMs and automated log templates will greatly improve log clustering, thereby enhancing coarse-grain debugging by efficiently grouping related failures. Additionally, extending VCDiag to support scenarios with multiple concurrent faulty modules through an iterative ’onion-peeling’ methodology will significantly broaden its practical applicability. This iterative refinement, however, would require developing strategies to accommodate intermediate design modifications seamlessly. Finally, focusing on improving the efficiency of regression reseeds, either through optimized simulation strategies or intelligent prioritization of test scenarios, will accelerate bug dataset generation and bring VCDiag closer to an industry-ready ML debugging framework.
@@ VII Conclusion
VIII Acknowledgment
This work was partially supported by the National Science Foundation (CCF-2106725) and the Hans Fischer Senior Fellowship by the Technical University of Munich Institute for Advanced Study.